OpenStudy (anonymous):

log36(6)=1/2 How do I write this in exponential form?

OpenStudy (jackellyn):

The logarithmic form log36 (6) = 1 / 2 is equivalent to the exponential form 6 = 36^(1/2)

OpenStudy (anonymous):

Thank you !

OpenStudy (jackellyn):

\[\log _{b} a = c\] in exponential form is b=a^c

OpenStudy (jackellyn):

No Problem